OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. 8 + 16 + 24 + ... + 8n = 4n(n + 1)

OpenStudy (anonymous):

@Shalante ?

ganeshie8 (ganeshie8):

Is the statement true for \(n=1\) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

how, can u explain

OpenStudy (anonymous):

4(1)(1+1) 4(2) =8

ganeshie8 (ganeshie8):

that means the expression \(4n(n+1)\) correctly gives the sum on left hand side for \(n=1\)

ganeshie8 (ganeshie8):

what if you have 2 terms on left hand side : \(8+16\) is the statement still true ?

OpenStudy (anonymous):

wait... you lost me. i have to plug in both now?

ganeshie8 (ganeshie8):

you don't need to, one base case is sufficient here, do you know how induction works ?

OpenStudy (anonymous):

i'm a little rusty on induction.

ganeshie8 (ganeshie8):

basically your goal is to prove that the given statement is true for all positive integers : \(n=1,2,3,\ldots \)

OpenStudy (anonymous):

i understand that. but I get confused after I've already proven 1

OpenStudy (anonymous):

i don't know where to go from there

ganeshie8 (ganeshie8):

suppose, you "know" that the given statement is true for \(n=k+1\), whenever it is true for \(n=k\)

ganeshie8 (ganeshie8):

is that sufficient to claim that the given statement is true for all integers : \(n=1,2,3,\ldots\) ?

OpenStudy (anonymous):

so then i have to plug in k+1

ganeshie8 (ganeshie8):

forget about the process, il surely give you the complete proof later for now, im just asking you a question..

ganeshie8 (ganeshie8):

Here it is again : suppose, you "know" that the given statement is true for \(n=k+1\), whenever it is true for \(n=k\). is that sufficient to claim that the given statement is true for all integers : \(n=1,2,3,\ldots\) ?

OpenStudy (anonymous):

no?

ganeshie8 (ganeshie8):

why no ?

OpenStudy (anonymous):

honestly idk

ganeshie8 (ganeshie8):

if you know that the statement is true for \(n=k+1\), whenever it is true for \(n=k\). that essentially means the statement is true for ALL positive integers if you could establish it for \(n=1\)

ganeshie8 (ganeshie8):

because, once you know that it is true for \(n=1\), it follows that it must be true for \(n=1+1=2\)

ganeshie8 (ganeshie8):

after you know that it is true for \(n=2\), it follows that it must be true for \(n=2+1=3\)

ganeshie8 (ganeshie8):

after you know that it is true for \(n=3\), it follows that it must be true for \(n=3+1=4\)

ganeshie8 (ganeshie8):

see the pattern ? it follows that it must be true for ALL integers greater than 1

OpenStudy (anonymous):

ok i understand that

ganeshie8 (ganeshie8):

So all we need to do is to establish two things : 1) the statement is true for \(n=1\) 2) the statement is true for \(n=k+1\), whenever it is true for \(n=k\)

ganeshie8 (ganeshie8):

then by the previous induction patter, it will follow that the statement is true for ALL integers greater than 1

ganeshie8 (ganeshie8):

You have already proven that the statement is true for \(n=1\) so lets look at the second step

OpenStudy (anonymous):

ok so what im doing is plugging in k+1?

ganeshie8 (ganeshie8):

Suppose the statement is true \(n=k\), then we have : \[8 + 16 + 24 + ... + 8k = 4k(k + 1)\]

ganeshie8 (ganeshie8):

add \(\color{red}{8(k+1)} \) both sides and get : \[8 + 16 + 24 + ... + 8k \color{red}{+ 8(k+1)} = 4k(k + 1)\color{red}{+ 8(k+1)} \]

ganeshie8 (ganeshie8):

factoring out 4(k+1) on right hand side, do we get : \[8 + 16 + 24 + ... + 8k \color{red}{+ 8(k+1)} = 4(k + 1)(k+2) \] ?

OpenStudy (anonymous):

i'm sorry but all those parenthesis kinda confuses me.... O.O

ganeshie8 (ganeshie8):

thats okay, take ur time

OpenStudy (anonymous):

i'm sorry but can you type it a bit more differently?.... is this what you're saying? 8+16+24+...+8k+8(k+1) = 4(k+1)(k+2)

ganeshie8 (ganeshie8):

Yes, is there any issue with latex ?

ganeshie8 (ganeshie8):

is it getting displayed like below on ur side ? https://i.gyazo.com/70406dea3fd312ded38a25e3f30da15b.png

OpenStudy (anonymous):

its coming out like this: \[8 + 16 + 24 + ... + 8k \color{red}{+ 8(k+1)} = 4(k + 1)(k+2) \]

ganeshie8 (ganeshie8):

take a screenshot and attach, browser messes up the latex sometimes

OpenStudy (anonymous):

how do i screenshot it?

ganeshie8 (ganeshie8):

https://gyazo.com/

OpenStudy (anonymous):

https://gyazo.com/e2eb27136301fb6cfe981d523de764f0

ganeshie8 (ganeshie8):

Ahh that looks horrible, I'll give you the screenshot of complete proof, one sec

ganeshie8 (ganeshie8):

Here is the complete proof https://i.gyazo.com/37b4e3e50a893db312121d7ed5de0944.png

ganeshie8 (ganeshie8):

let me knw if something doesn't make sense

OpenStudy (anonymous):

oh ok i see.

OpenStudy (anonymous):

i understand it. thank you alot.

ganeshie8 (ganeshie8):

yw